Status Report

NASA Hubble Space Telescope Daily Report # 3802

By SpaceRef Editor
February 22, 2005
Filed under , ,
NASA Hubble Space Telescope Daily Report # 3802

HUBBLE SPACE TELESCOPE – Continuing to collect World Class Science

DAILY REPORT        # 3802




ACS & WFPC2 Two-gyro PSF, pointing and dither test

12 orbits High priority observations {5 orbits} will be observed
together at start of two- gyro SMS as contingency against HST failure.
Remaining 7 orbits provide further tests and additional experience in
two-gyro mode. It is believed that this will be the only test prior to
final failure into two-gyro mode. Goals: PSF, PSF repeatability,
pointing and stability within the orbit {requirements 1a, 1c, 2 , 4}

FGS 10457

2-Gyro F2G Test Miscellaneous Orbits

This proposal is part of the February on-orbit 2-gyro testing.

NIC1/NIC2 10447

NICMOS Dither Test two-gyro mode

The accuracy of the combined dither and chop mode in NICMOS running in
two-gyro mode is tested. Dithering permits the removal of dead or
non-calibrated pixels on the detectors and improves the spatial
sampling and the effects of detector non- uniformities. Chopping is
used for mapping and background subtraction. The latter can involve
large offsets in order to find blank sky. In this test we use the
NIC-SPIRAL- DITH-CHOP pattern with a dither step of 1.5″ and a chop of
100″. The chop distance is small enough that a re-acquisition of the
guide star is not required {chosen before larger chop distance because
of time constraint}. NIC1 and NIC2 cameras are involved, where NIC2 is
primary and NIC1 parallel.

NIC1/NIC2/NIC3 10446

NICMOS 2Gyro PSF-test

This proposal is aimed at analyzing the NICMOS PSF under 2-gyro
operations. For that we will observe the standard star P330E, which
has been intensively observed since Cyle 7 and, therefore, it is a
good reference to detect any possible deviation from 3- gyro
operations. This proposal is largely based on the photometric
monitoring calibration proposal {10381 and previous}. Changes with
respect to that proposal include: 1- The patterns are repeated twice
for NIC1, and 3 times for NIC2 in order to gain S/N and reduce
intrapixel sensitivity. NIC3 observations {1 repeat} complete the last
orbit. 2- For F110W observations the SCAMRR sequence is used instead
of MCAMRR, in order to gain temporal resolution. 3- Filters F190N and
F222M will not be observed with NIC3. The effects of jitter on these
filters should be very small {and better detected with NIC2}.
Intrapixel sensitivity variations induced by the jitter should be
better studied at shorter wavelengths.

NIC1/NIC2/NIC3 8792

NICMOS Post-SAA calibration – CR Persistence Part 3

A new procedure proposed to alleviate the CR-persistence problem of
NICMOS. Dark frames will be obtained immediately upon exiting the SAA
contour 23, and every time a NICMOS exposure is scheduled within 50
minutes of coming out of the SAA. The darks will be obtained in
parallel in all three NICMOS Cameras. The POST-SAA darks will be
non-standard reference files available to users with a USEAFTER
date/time mark. The keyword ‘USEAFTER=date/time’ will also be added to
the header of each POST-SAA DARK frame. The keyword must be populated
with the time, in addition to the date, because HST crosses the SAA ~8
times per day so each POST-SAA DARK will need to have the appropriate
time specified, for users to identify the ones they need. Both the raw
and processed images will be archived as POST-SAA DARKSs. Generally we
expect that all NICMOS science/calibration observations started within
50 minutes of leaving an SAA will need such maps to remove the CR
persistence from the science images. Each observation will need its
own CRMAP, as different SAA passages leave different imprints on the
NICMOS detectors.

NIC2 10448

NICMOS 2-gyro Coronagraphic Performance Assessment

Coronagraphic acquisition should not be affected by transitioning from
3-gyro to 2-gyro mode guiding. However, coronagraphic performance is
dependent upon positioning of a target within the coronagraphic hole
and the ability to obtain a second observation with a roll of the
telescope within the same visibility period. Observations of a bright
target are needed to verify and confirm the coronagraphic performance,
the ability to suppress the diffractive energy background and the
stability to image coronagraphically occulted targets at two field
orientations {rolls}.


Significant Spacecraft Anomalies: (The following are preliminary
reports of potential non-nominal performance that will be

HSTAR 9714 : GSacq(1,3,1) failed to RGA Control @ 052/1331z – The
GSAcq(1,3,1) scheduled at 052/13:31:32 achived F2G-FL, then
transitioned to RGA control (1.e) F2G-FL/T2G/M2G Mode. Stop flag
indication was observed. Prior OBAD at 052/13:11:45, and 052/13:23:12
were successful.  Possible Observations affected:  NIC 11 thru 16.  
Under investigation.

HSTAR 9716 : GSAcq(1,3,1) failed to RGA control during F2G On-Orbit
Test @ 052/1507z – The GSAcq(1,3,1) scheduled at 052/15:07:32 during
F2G On-Orbit Test achived F2F-FL, then failed to RGA control (i.e)  
transitioned to F2G/T2G Mode. Stop indication flag was observed. Same
guide star id per ref HSTAR 9714.  Prior OBAD at 052/14:47:42, and
052/14:59:12 were successful.Possible Observations affected: NIC 17
thru 22. Under investigation.

17370-0 – Continuous Engineering Coverage for F2G OOT (ESTR2 Power ON) @

17370-0 – Continuous Engineering Coverage for F2G OOT (ESTR2 Record to 18.41%)
                @ 052/1440z-1535z

17383-0 – Patch Acq. Logic for Target ReAcq @ 053/0017z

17370-0 – Continuous Engineering Coverage for F2G OOT (ESTR2 Record  to
                @ 053/0117z-0210z

17370-0 – Continuous Engineering Coverage for F2G OOT (ESTR2 to BOT) @


                              SCHEDULED     SUCCESSFUL    FAILURE TIMES 
 FGS Gsacq                   7                         7 
 Reacq                   7                         5               052/1331z 
 (HSTAR 9714) & 
 (HSTAR 9716) 
 FHST Update 
 OBAD                           28                      28 


The 3-day Two-Gyro Science (TGS) on-orbit test is in progress.

This test was the first use of a STScI generated TGS science mission
timeline (SMS). HST was commanded into Two-Gyro Mode on Feb 20th at
~8:00 pm EST.  The new Fine Guiding Mode (F2G) mode is performing
well, and HST boresight jitter is similar to PCS predictions.

SpaceRef staff editor.